OpenStudy (anonymous):

1/ tanx + cosx = tanx = sinx prove the identity

OpenStudy (turingtest):

got a typo there...

OpenStudy (turingtest):

1/ tanx + cosx = tanx + sinx maybe?

OpenStudy (anonymous):

yeah sorry

OpenStudy (anonymous):

its 1/ tanx + cosx = tanx + sinx

OpenStudy (turingtest):

it's not true... http://www.wolframalpha.com/input/?i=1%2Ftanx%2Bcosx%3Dtanx%2Bsinx it's not an identity...

OpenStudy (ash2326):

yeah it's not true, check your question again

OpenStudy (anonymous):

how is it not an identity? thats the right answer apparently

OpenStudy (anonymous):

ash can u answer my question above this post

OpenStudy (ash2326):

this question? it's a wrong identity

OpenStudy (anonymous):

the one that says 13 minutes ago

OpenStudy (ash2326):

where is it?

OpenStudy (anonymous):

below your first post in this question

OpenStudy (ash2326):

OK, i checked it, there must be an error to prove 1/tan x+ cos x= sin x+tanx

OpenStudy (anonymous):

how is it not an identity? thats the right answer apparently

OpenStudy (ash2326):

1/tan x+ cos x* tan x / tan x (1+sin x)/ tan x this is the end we can also prove this wrong by replacing x with a value say x=pi/6 sin pi/6= 1/2 cos pi/6=root 3/2 tan pi/6= 1/root 3 LHS we have \[1/\tan (\pi/6)+ \cos \pi/6\] \[\sqrt 3+ \sqrt 3/2\] RHS we have tan pi/6 + sin pi/6 \[1/ \sqrt 3+ 1./2\] so LHS is not equal to RHS, this identity is not valid

OpenStudy (anonymous):

ok thanks